In my absence bringing a van back from Surrey yesterday, Ian reports a bright sunny but cold morning at Grosmont with Jon Bradley, Bill Dobson, Steve Hyman, Derek Shorten, Nigel Hall, and Ian himself. Later on in the morning Roy Marshall called in briefly, and feeling a bit under the weather, to deliver a couple of new strops and some paperwork for Bryan Orange (which I collected today and will give to Bryan on Saturday). No mice again, but two traps are now set!
Usual cups of tea all round that cold morning to warm everyone up, then work started with Steve removing the inner part of the valve gland follower, as it had moved slightly out of line with the outer part causing the oil supply to be cut off and the valve spindle not getting lubricated. But just as he had got it apart, a call came from the MPD asking if we could spare a couple of our team to help Andrew Jeffery and Jamie Lazenby down at the MPD on No 5 road to weigh the J27 tender. So Steve and Jon spent three parts of the day at the MPD. After completing the weighing Andrew gave us a copy of the weighing results: these are in the red In box on the table upstairs in Deviation Shed (photo shows the weighing results on the electronic display and the tender being weighed). In the meantime, Bill spent some time at the MPD trying to see where we were in the queue with various jobs, from boiler work to lathe work. At the moment we still seem to be at the back of the queue, although Mark has now welded the nicks on the front tube plate. Nigel continued to repair the valve follower left by Steve. The inner and outer sections were lined up, drilled and secured with a small grommet - job complete and ready for fitting. Derek finished painting the tender steam heat and vacuum pipes then set about cleaning the grime and grease from the tender wheels. Ian continued working underneath the locomotive, fitting taper pins on the eccentric rods.
Lunch was had in Deviation shed seeing as the weather was a little better, not too bad for us hardy Northern souls! Later, Jon oiled and put the trimmings into the eccentric oil reservoirs and new corks were fitted. Bill drilled through the small boss from the trailing R/H crankpin, which now requires a small taper reamer to complete the job. That will be completed on Saturday. Roy brought 2 new, 1 metre, 1 ton strops, Nos 510190 & 510186. These are now with the others in the workshop and are OK to use. Nigel found one old strop hanging from the Q6 dome - 1 metre, 1 ton, No. 880172 - which was missed at the last inspection. This looks to be in good order but cannot be used until the next inspection in six months time. It is now in the filing cabinet with the other two that were found a few weeks ago: we must try and put this equipment back in the workshop after use instead of being left out and forgotten about! All the paper records are in a red file in the workshop.
Standard Class 4 No 76079 and Class 25 diesel were working top and tail on Esk Valley Line route refreshers. There was also lots of shunting round the MPD moving the J27 out so Eric Treacy could be put outside ready for its boiler lift today. The B1 and Standard Tank No 80136 were on steam tests.
Next working party will be on Saturday 24th March. Plenty of milk and biscuits in stock. Remember though that Saturday sees the start of the 2018 operating season, albeit with the winter timetable used that day, Silver on Sunday and then into Red for the week before and after Easter. That means the station cafe will be open, but also that the car park will be busier than we are used to, and there may be rather more visitors about - there were certainly a lot there today when I was briefly at Grosmont checking on progress with the J27, for which a separate note will follow in due course.